OpenStudy (anonymous):

PLEASE HELP!!!!! the equation of a conic section is 2x²-10y²-12x-2=0 , what is the equation in graphing form?

OpenStudy (amistre64):

not too sure what is meant by "graphing form" but the setup looks hyperbolic to me, try completeing the squares for the x and y parts

OpenStudy (amistre64):

2x²-10y²-12x-2=0 2x²-10y²-12x = 2 x²-5y²-6x = 1 shuffling tings about might give you a feel of a simpler setup